### v1

# Marked length spectrum rigidity of group acts isometrically on a Gromov hyperbolic space

Abstract.
We generalize the result of Nguyen and Wang: given a relatively hyperbolic group with two left-invariant metrics quasi-isometric to the cone-off word metric on it, if the marked length spectrum (the translation length function) under the two metrics are the same, then the two metrics are uniformly close, to the case of certain subgroups of the isometric group of a proper Gromov hyperbolic spaces. We also give a sparse, spectrally rigid subset with at most two limit points, which is connected to a series of works by Cantrell and Butt.
Key words and phrases:
marked length variety, marked length pattern, arithmetic surface
2020 Mathematics Subject Classification:
37D40; 37A20
1. Introduction
Given a closed Riemannian manifold $(M,g)$ with negative sectional curvatures and its fundamental group $\pi_{1}(M)=\Gamma$, a unique closed geodesic exists within each free homotopy class representing a non-trivial conjugacy class of $\Gamma$. Denoting the set of these conjugacy classes of $\Gamma$ as $\mathcal{C}$, we define the marked length spectrum map for $g$:
$\ell_{g}:\mathcal{C}\to\mathbb{R}^{+}$
by associating each class with the length of the corresponding closed geodesic.
A coarsely geometric generalization of the marked length spectrum map is also applicable. Consider a Gromov hyperbolic space $(X,d)$ along with an isometry $f:X\to X$ an isometry, and a fixed base point $o\in X$. We define the translation length of $f$ as
$\ell_{d}(f)=\lim_{n\to\infty}\frac{d(f^{n}(o),o)}{n}.$
This definition is independent of the choice of the base point $o\in X$. When $(X,d)$ serves as the Riemannian universal cover of $(M,g)$, and $f\in\Gamma=\pi_{1}(M)$ acts as a deck transformation, the stable length coincides with the marked length spectrum map. Consequently, we refer to the map $\ell_{d}$ as the marked length spectrum of the isometry group of $X$.
In this context, Furman [F] established the following rigidity theorem:
Theorem 1.1.
Consider a countable hyperbolic group $\Gamma$, and let $d_{1}$ and $d_{2}$ be two roughly geodesic left invariant metrics on $\Gamma$, which are quasi-isometric to a word metric (with a finite set of generators). If $\ell_{d_{1}}=\ell_{d_{2}}$ on $\Gamma$, i.e., for all $\gamma\in\Gamma$, $\ell_{d_{1}}(\gamma)=\ell_{d_{2}}(\gamma)$, then there exist a constant $C$ such that $|d_{1}-d_{2}|<C$.
Fujiwara [fujiwara2015asymptotically], built upon a strategy initially proposed by Krat [krat2001pairs], extended this result to toral relatively hyperbolic groups. More recently, Nguyen and Wang [nguyen2022marked] further confirmed this result for all relative hyperbolic groups. It’s important to note that all these results pertain to situations where the metrics are quasi-isometric to a word metric. However, we now raise the question: what about the general case? We try to answer this question in a simple case: the case of isometric actions on Gromov hyperbolic spaces.
Our initial focus is on negatively curved non-compact manifolds with finite volume and bounded geometry. We present the following theorem:
Theorem A.
Let $M$ be a manifold of dimension $d\geq 2$. Suppose $g_{1}$ and $g_{2}$ are two complete, finite volume, Riemannian metrics on $M$ with sectional curvature satisfying $-a^{2}\leq\kappa\leq-b^{2}<0$. If the marked length spectrum of $g_{1}$ and $g_{2}$ coincide, then the universal covers of the two metrics are roughly isometric.
On the other hand, there’s a genuine desire to comprehend the partial rigidity of the marked length spectrum. The study of Riemannian manifolds in this context has a rich history. It is often the case that information about a proper yet well-structured subset is sufficient to determine the entire marked length spectrum. Katok [katok1988four] showed that given a closed manifold $M$, for any homology class $[c]\in H_{1}(M,\mathbb{Z})$, two negatively curved metrics in a fixed conformal class possess the same marked length spectrum if and only if the marked length spectrum map is in agreement for all loops representing $[c]$. Gogolev and Rodriguez Hertz [gogolev2020abelian] strengthened Katok’s result by showing that it still holds even when the condition ‘in a fixed conformal class’ is dropped. See Remark 6.5 in [gogolev2020abelian]. In our previous work [hao2022marked], we showed that it is enough to consider the restriction of the marked length spectrum to a subgroup with the same limit set as the fundamental group.
One of the principal focuses of this paper is to extend these results to certain groups that act isometrically on a ‘good’ Gromov hyperbolic space.
Before delving further, let us establish some notation. Consider $(X,d)$, a $K$-roughly geodesic, $\delta$-Gromov hyperbolic space. We call $X$ $K$-visual if for all points $p\in X$ and $x\in\partial X$, there exists a $K$-roughly geodesic start from $p$ and converge to $x$. We label $(X,d)$ as $K$-roughly complete if, for any two point $p$, $q\in X$, there exists an infinite $K$-rough geodesic path passing $p$, $q$ and converging to some $x\in\partial X.$
Definition 1.2.
A metric space $(X,d)$ is called a roughly hyperbolic manifold if there exist $K$, $\delta>0$ such that the following are satisfied:
  1. (1)
    $(X,d)$ is $K$-roughly geodesic, $\delta$-Gromov hyperbolic space;
  2. (2)
    $(X,d)$ is $K$-visual;
  3. (3)
    $(X,d)$ is $K$-roughly complete.
We call a hyperbolic manifold nontrivial if the Gromov boundary $\partial X$ contains at least 3 points.
Example 1.3.
Let $\Gamma$ be a hyperbolic group. Any word metric with finitely many generators is a nontrivial roughly hyperbolic manifold.
Given a Gromov hyperbolic space $(X,d)$, a non-elementary subgroup $\Gamma$ of $\operatorname{Isom}(X,d)$ is dense at infinity if the limit set of $\Gamma$, $\Lambda\Gamma$ is the same as $\partial X$, the boundary of $X$. Given two actions of $\Gamma$ on $(X,d)$ and $(Y.d^{\prime})$, respectively, which are dense at infinity, these two actions are called comparable if a $\Gamma$-equivariant, continuous homeomorphism exists from $\partial X$ to $\partial Y$.
In this paper, we only consider actions that are dense at infinity.
There are many examples of actions dense at infinity. We give two that is our may concern.
Example 1.4.
A normal (or commensurated) subgroup of a hyperbolic group acts on the hyperbolic group with word metric.
Example 1.5.
Consider a dense (in Hausdorff topology) finite rank free subgroup $F$ of $\operatorname{PSL}_{2}(\mathbb{R})$. The $F$-action on the hyperbolic plane $H^{2}$ is dense at infinity.
The following is the first main theorem of this paper.
Theorem B.
Let $\Gamma$ be a group. Given two comparable actions of $\Gamma$ on nontrivial roughly hyperbolic manifolds $(X,d)$ and $(Y,d^{\prime})$, if $\ell_{d}=\ell_{d^{\prime}}:\Gamma\to\mathbb{R}^{\geq 0}$, then there exist a $\Gamma$-equivariant roughly isometry $\phi:X\to Y$, i.e., there exist a constant $C>0$, such that for all $p$, $q\in X$,
$|d^{\prime}(\phi(p),\phi(q))-d(p,q)|<C.$
Remark 1.6.
It is clear that Theorem A is a consequence of Theorem B. Hence, we will only prove the latter.
The proof strategy is similar to Furman’s work [F]. First, in Furman’s work, $X$ and $Y$ have the marked length spectrum, implying that their cross-ratios on the boundary are the same. We define rough cross-ratio on the boundary and show that the marked length spectrum roughly determines the rough cross-ratio. The advantage of the rough cross-ratio is that unlike Furman’s definition, which is merely a measurable function, the rough cross-ratio is defined everywhere, though not continuous. Then, we consider the space of metrics $\mathcal{D}^{R}$ on the boundary with the same rough cross-ratio and bounded geometry. Applying a generalization of the barycenter method used in [biswas2015moebius], $X$ and $Y$ are roughly isometric to $\mathcal{D}^{R}$. Therefore, Theorem B follows.
The space we constructed in the second step is a hyperbolic filling of the boundary. There are also other ways of constructing hyperbolic filling. One standard method is to compare hyperbolic spaces of constant curvature. Previous work in this direction includes Paulin [paulin1996groupe], Bonk and Schramm [bonk2011embeddings], Buyalo and Schroeder [buyalo2007elements], and Jordi [jordi2010interplay], among others. It is possible to use these constructions for our purpose. However, we apply the barycenter method since it is closely related to step one in our proof. See section 2.D for a more detailed discussion.
Note that in Katok [katok1988four] and Gogolev and Rodriguez Hertz [gogolev2020abelian], the subsets considered are not a subgroup in general. We also have a generalization for coset of a subgroup in this form.
Theorem C.
Let $\Gamma$ be a group. Given two comparable actions of $\Gamma$ on roughly hyperbolic manifolds $(X,d)$ and $(Y,d^{\prime})$. Assume there exist a subgroup Γsuperscriptnormal-$\Gamma^{\prime}$ and an element $h\in\Gamma$ so that the limit set $\Lambda\Gamma^{\prime}=\Lambda\Gamma$ for both actions, and $\ell_{d}=\ell_{d^{\prime}}$ on the left coset $h\Gamma^{\prime}$ or the right coset $\Gamma^{\prime}h$, then $\ell_{d}=\ell_{d^{\prime}}$ on $\Gamma$, and there exist a $\Gamma$-equivariant roughly isometry $\phi:X\to Y$, i.e., there exist a constant $C>0$, such that for all $p$, $q\in X$,
$|d^{\prime}(\phi(p),\phi(q))-d(p,q)|<C.$
Remark 1.7.
Since the two actions are comparable, the condition $\Lambda\Gamma^{\prime}=\Lambda\Gamma$ for one action is enough.
Remark 1.8.
Note that it is essential for the space the group acts on to be roughly geodesic since, in general, a $\Gamma$-orbit is not roughly geodesic as a metric space. An example will be a non-uniform lattice $\Gamma$ of $\operatorname{PSL}_{2}(\mathbb{R})$ acts on the hyperbolic plane $H^{2}$. Any $\Gamma$ orbit is not roughly geodesic with the induced metric. On the other hand, for Theorem D and Theorem E in the next subsection, it is enough to consider only one $\Gamma$-orbit.
1.A. Spectrally rigid set
In exploring additional avenues of marked length spectrum rigidity, we delve into spectrally rigid sets. Consider a group $\Gamma$ and two comparable actions of $\Gamma$ on roughly hyperbolic manifolds $(X,d)$ and $(Y,d^{\prime})$. As in [cantrell2023sparse], we say a collection of conjugacy class $E$ spectrally rigid if, when $l_{d}([x])=\ell_{d^{\prime}}([x])$ for all $[x]\in E$, it follows that $\ell_{d}=\ell_{d^{\prime}}$ for all $[x]\in\mathcal{C}$. Sawyer [sawyer2020partial] has shown that the coinciding marked length spectrum on a set, whose complement exhibits sub-exponential growth, uniquely determines the marked length spectrum of a compact surface. Cantrell and Cantrell-Reyes, in a series of papers [cantrell2022manhattan, cantrell2023marked, cantrell2023sparse], extended this notion to a coarse-geometric setting, yielding remarkable results. Notably, Cantrell constructed arbitrarily sparse spectrally rigid sets for hyperbolic groups.
We have a similar result here.
Theorem D.
Let $f:\mathbb{R}\to\mathbb{R}^{+}$ such that $\lim_{T\to\infty}f(T)=\infty$, and $\Gamma$ be a group. There exists a subset $E\subset\Gamma$ such that, given two comparable actions of $\Gamma$ on two non-trivial Gromov hyperbolic spaces $(X,d)$ and $(Y,d^{\prime})$ that are dense at infinity, if $\ell_{d}(\gamma)=\ell_{d^{\prime}}(\gamma)$ for all $\gamma\in E$, then $\ell_{d}=\ell_{d^{\prime}}$ on $\Gamma$. Furthermore, we can choose $E$ such that:
  1. (1)
    the growth of $E$ with respect to $d$ (or/and $d^{\prime}$) satisfies:
    #$\#\{\gamma\in E|\ell_{d}(\gamma)\leq T\}\leq f(T).$
  2. (2)
    If the action is quasi-parabolic, for ant $\gamma\in\Gamma$ acts on $X$ hyperbolically, the limit set of $E$ contains only the $\gamma$-fixed point which is not $\Gamma$-fixed.
  3. (3)
    If the action is non-elementary, for any two given $\gamma$ and $\theta\in\Gamma$ with $\gamma$ acts on $X$ hyperbolically and $\theta\gamma^{-}\neq\gamma^{-}$, the limit set of $E$ contains at most two points: $\gamma^{-}$ and $\theta\gamma^{-}$.
While the methodology in [cantrell2023sparse] is expected to be applicable for cofinite volume actions, certain challenges arise. Firstly, as noted in [cantrell2023marked]*page 2, last paragraph, ‘geodesic flow and Busemann cocycle are in general not available outside the Riemannian or low dimensional setting.’ In this paper, we introduce a rough version of the ‘Busemann cocycle,’ a valuable tool in numerous cases. Secondly, the generalization of their work to find spectrally rigid subsets of general subgroups of $\operatorname{Isom}(X)$ remains uncertain. Hence, we propose an alternative construction that yields an arbitrarily sparse spectrally rigid set.
The foundational work of Butt [butt2022quantitative, butt2022approximate] serves as the inspiration for the last ingredient of this paper. Butt ingeniously proposed that for two distinct Riemannian metrics of negative curvature on a closed manifold of dimension at least 3, a closely matched marked length spectrum implies their volumes are closely aligned. Additionally, she showed that, under certain geometric conditions, the values of the marked length spectrum on a sufficiently large finite set can effectively determine the metric. In this work, we extend this result to encompass all actions on Gromov hyperbolic spaces that are dense at infinity.
Theorem E.
Let $f:\mathbb{R}\to\mathbb{R}^{+}$ such that $\lim_{T\to\infty}f(T)=\infty$ and $\Gamma$ be a torsion-free hyperbolic group. There exists a subset $E\subset\Gamma$ consisting of primary elements such that, for two comparable faithful actions of $\Gamma$ on two nontrivial Gromov hyperbolic spaces $(X,d)$ and $(Y,d^{\prime})$, if there is $C>1$ with $\frac{1}{C}\ell_{d^{\prime}}(\gamma)\leq\ell_{d}(\gamma)\leq C\ell_{d^{\prime}}(\gamma)$ for all $\gamma\in E$, then $\frac{1}{C}\ell_{d^{\prime}}(\gamma)\leq\ell_{d}(\gamma)\leq C\ell_{d^{\prime}}(\gamma)$ for all $\gamma\in\Gamma$. Furthermore, we can choose $E$ so that the growth of $E$ with respect to $d$ (or/and $d^{\prime}$) satisfying:
#$\#\{\gamma\in E|\ell_{d}(\gamma)\leq T\}\leq f(T).$
Remark 1.9.
It is crucial that $E$ only contains primary elements. For otherwise, taking $E^{\prime}=\cup\{\gamma^{n_{\gamma}}\}$ over all primary conjugacy class for some sequence of integers $n_{\gamma}$, $E^{\prime}$ would determines the marked length spectrum of $\Gamma$. By choosing $n_{\gamma}$ big enough, $E$ can be made arbitrarily sparse.
We term a subset $E$ detecting the ratio of the marked length spectrum if it satisfies Theorem E.
The author expresses gratitude to Ralf Spatizer for valuable comments and for drawing attention to the works of Butt and Cantrell. Thanks to Cantrell for many suggestions and comments. Additionally, thanks to Ashot Minasyan for highlighting reference [minasyan2010normal].
2. Roughly geometric hyperbolic space
2.A. Gromov product and Gromov hyperbolic spaces
Some good general references for Gromov hyperbolic spaces are [gromov1987hyperbolic] and [bridson2013metric]*Chapter III.H Section 1 and 3. We recall some basic definitions and results here.
Let $(X,d)$ be a metric space with a base point $o$. The Gromov product $<,>_{o}:X\times X\to\mathbb{R}^{\geq 0}$ is given by
$<p,q>_{o}=\frac{1}{2}(d(p,o)+d(q,o)-d(p,q)).$
Let $\delta>0$. $(X,d)$ is called $\delta$-Gromov hyperbolic if for all $o$, $p$, $q$, $r\in X$,
$<p,q>_{o}\geq\min\{<p,r>_{o},<r,q>_{o}\}-\delta.$
Note that if there exists some $o\in X$ such that the above inequality holds for all $p$, $q$, $r\in X$, then $X$ is $2\delta$-Gromov hyperbolic. See [gromov1987hyperbolic]*1.1 B. We say $(X,d)$ a Gromov hyperbolic space if it is $\delta$-Gromov hyperbolic space for some $\delta$.
Let $(X,d)$ be a hyperbolic space. A sequence $(x_{n})$ in $X$ converges at infinity if $<x_{i},x_{j}>_{o}\to\infty$ as $i$, $j\to\infty$. Two such sequence $(x_{n})$ and $(y_{n})$ are said to be equivalent if $<x_{i},y_{j}>_{o}\to\infty$ as $i$, $j\to\infty$. The equivalence class of $(x_{n})$ is denoted $\lim x_{n}$, and the set of equivalence classes is denoted $\partial X$: the Gromov boundary of $X$.
The Gromov product of a $\delta$-hyperbolic space $X$ can be extended to $\bar{X}=X\cup\partial X$ by:
$<x,y>_{o}=\sup\liminf_{i,j\to\infty}<x_{i},y_{j}>,$
where the supremum is taken over all sequences $(x_{i})$ and $(y_{j})$ in $X$ such that $x=\lim x_{i}$ and $y=\lim y_{j}$.
Remark 2.1.
[…]
(2.1) $<x,y>_{o}-2\delta\leq\liminf_{i,j}<x_{i}^{\prime},y_{j}^{\prime}>_{o}\leq<x,y>_{o}.$
See [bridson2013metric]*pp 432-433.
Note that the extended Gromov product is not continuous in general. Nevertheless, there is a rough continuity. Recall (see [bridson2013metric]*pp 433) that for all $x$, $y$, $z\in\bar{X}$, we have
(2.2) $<x,y>_{o}\geq\min\{<x,z>_{o},<x,z>_{o}\}-2\delta,$
and the topology on $\bar{X}$ is given by $x_{n}\to x$ if $<x_{n},x>_{o}=\infty$ or $x_{n}$ converge to $y$ in $X$, it follows for any sequence $(y_{n})$ in $\bar{X}$ with $y_{n}\to y$ and $z\in\bar{X}$, we have
(2.3) $|\liminf<y_{n},z>-<y,z>|\leq 2\delta.$
When $y\in X$, this follows by taking the limit of the triangle inequality.
In the last part, we define the (logarithmic) rough cross-ratio. Let $\partial^{(4)}X$ be the set of pairwise different 4-tuples in $\partial X$. The rough cross-ratio is defined by the map
(2.4) $(x,y;z,w)=<x,y>_{o}+<z,w>_{o}-<z,y>_{o}-<x,w>_{o}.$
Note that when $(X,d)$ is a CAT(-1) space, our definition is proportional to the logarithm of the standard cross-ratio defined by Otal [otal1992geometrie].
2.B. Quasi-isometries, roughly geodesic space
We follow the definition in [bonk2011embeddings]. A map $f:(X,d)\to(Y,d^{\prime})$ of metric spaces is call a $(K,C)$-quasi-isometric embedding if
$\frac{1}{K}d^{\prime}(f(p),f(q))-C\leq d(p,q)\leq Kd^{\prime}(f(p),f(q))+C,$
for all $p$, $q\in X$. A $(K,C)$-quasi-isometric embedding is called a quasi-isometry if the image is cobounded, i.e., there exist constant $L$ such that for all point $r\in Y$, there exist $p\in X$, with $d^{\prime}(r,f(p))<L$.
$(1,C)$-quasi-isometric embedding and quasi-isometric are called $C$-roughly isometric embedding and $C$-rough isometry. A $C$-rough geodesic is a $C$-roughly isometric embedding of an interval. A metric space is called $C$-roughly geodesic if any two points can be connected by a $C$-rough geodesic.
Let $\tau(t)$ be a rough geodesic, for any $t_{1}<t_{2}<t_{3}$, we have
(2.5) $<\tau(t_{1}),\tau(t_{3})>_{\tau(t_{2})}\leq\frac{3}{2}C,$
(2.6) $<\tau(t_{2}),\tau(t_{3})>_{\tau(t_{1})}\geq|t_{2}-t_{1}|-\frac{3}{2}C.$
A rough geodesic space is called roughly complete if every rough geodesic extends to a bi-infinite rough geodesic.
2.C. Busemann cocycle and (rough) Busemann cocycle
Since its initial definition in [busemann2012geometry], the Busemann function has found applications in various problem domains. Notably, several applications involve its related counterpart, the Busemann cocycle, as discussed in [benoist2016central, nica2013proper, nica2019two]. We first recall the definition of the Busemann function. Let $(X,d)$ be a Gromov-hyperbolic space. For points $o,p\in X$ and $x\in\partial X$, the function $B:X\times X\times\partial X\rightarrow{\mathbf{R}}$ is defined by:
(2.7) $B_{o,p}(x)=\sup\liminf_{q_{i}\rightarrow x}d(o,q_{i})-d(p,q_{i})=2<p,x>_{0}-d(o,p),$
where $(q_{i})$ runs all the sequence converge to $x$.
Definition 2.2.
Let $(X,o,d)$ be a pointed Gromov hyperbolic space. The (rough) Busemann cocycle is the function:
$c:\operatorname{Isom}(X)\times\partial X\to\mathbb{R},$
$c(\gamma,x)=B_{o,\gamma o}(x).$
The extended Busemann cocycle is a similar function:
$c:\operatorname{Isom}(X)\times X\to\mathbb{R},$
$c(\gamma,p)=d(o,p)-d(\gamma o,p)=2<\gamma o,p>_{o}-d(o,\gamma o).$
Remark 2.3.
In earlier works, the Busemann cocycle is defined as the Radon-Nikodym derivative of the Patterson-Sullivan measure, constituting a cocycle. However, in that context, they are merely measurable functions, making it challenging to compare different Busemann functions without information about their Patterson-Sullivan measures. The definition adopted here maintains close to the Radon-Nikodym derivative, resulting in a rough cocycle that shares numerous properties. Importantly, the new functions are defined everywhere.
Here, we enumerate some properties of Busemann cocycles, with concise proofs provided for the reader’s convenience. Most of these properties are well-established, particularly when the Busemann function is continuous [nica2019two].
Proposition 2.4.
The Busemann cocycle $c$ is almost continuous. For any sequence $(y_{i})$ in $\bar{X}$ converging to $x\in\partial X$, and for any $\gamma\in\operatorname{Isom}(X)$,
$|\liminf_{i\to\infty}c(\gamma,y_{i})-c(\gamma,x)|\leq 4\delta.$
Proof.
This follows f directly from the definition of the Busemann function and equation (2.3). ∎
Proposition 2.5.
For any $\gamma$, $\gamma^{\prime}\in\operatorname{Isom}(X)$ and all $x\in\partial X$, we have
$|c(\gamma\gamma^{\prime},x)-c(\gamma,\gamma^{\prime}x)-c(\gamma^{\prime},x)|\leq 8\delta.$
Proof.
Choose $\epsilon>0$. Select a sequence $(q_{i})$ in $X$ converging to $x$ such that
$\lim_{i\to\infty}c(\gamma\gamma^{\prime},q_{i})-\epsilon\leq c(\gamma\gamma^{\prime},x)\leq\lim_{i\to\infty}c(\gamma\gamma^{\prime},q_{i})+\epsilon.$
Note that $c(\gamma\gamma^{\prime},q_{i})=c(\gamma,\gamma^{\prime}q_{i})+c(\gamma^{\prime},q_{i})$ for all $i$. By proposition 2.4],
$|\lim_{i\to\infty}c(\gamma,\gamma^{\prime}q_{i})-c(\gamma,\gamma^{\prime}x)|\leq 4\delta,$
$|\lim_{i\to\infty}c(\gamma^{\prime},q_{i})-c(\gamma^{\prime},x)|\leq 4\delta.$
Since $\epsilon$ is arbitrary, the result follows. ∎
A similar argument (setting proper sequence converge to $x$ and $y$) leads to:
Proposition 2.6.
For all $\gamma\in\operatorname{Isom}(X)$ and $x$, $y\in\partial X$,
$|<x,y>-<\gamma x,\gamma y>-c(\gamma,x)-c(\gamma,y)|\leq 4\delta.$
Proposition 2.7.
For all hyperbolic element $\gamma\in\operatorname{Isom}(X)$, $c(\gamma,\gamma^{\pm})=\pm\ell_{d}(\gamma)$.
Proof.
First, consider $c(\gamma,\gamma^{+})$. By the fact $\gamma\gamma^{+}=\gamma^{+}$ and Proposition 2.5, we have:
(2.8) $|c(\gamma^{n},\gamma^{+})-nc(\gamma,\gamma^{+})|\leq 8(n-1)\delta.$
Now, take any sequence $(q_{i})$ in $X$ converging to $\gamma^{+}$. Then, by proposition 2.4,
$|\liminf_{i\to\infty}c(\gamma^{n},q_{i})-c(\gamma^{n},\gamma^{+})|\leq 4\delta.$
Notice that
$\begin{array}[]{llr}c(\gamma^{n},q_{i})&=d(o,q_{i})-d(\gamma^{n}o,q_{i})&\text{(Definition)}\\ &=d(o,\gamma^{n}o)-2<o,q_{i}>_{\gamma^{n}o}&\\ &=d(o,\gamma^{n}o)-2<\gamma^{-n}o,\gamma^{-n}q_{i}>_{o}.&(\gamma\in\operatorname{Isom}(X))\end{array}$
Hence by equation (2.3)
$|\liminf_{i\to\infty}c(\gamma^{n},q_{i})-d(o,\gamma^{n}o)-2<\gamma^{-n}o,\gamma^{+}>_{0}|\leq 4\delta.$
Since $\gamma^{-n}o$ converge to $\gamma^{-}\neq\gamma^{+}$, we see that $<\gamma^{-n}o,\gamma^{+}>_{o}$ is bounded. Therefore,
$\lim_{n\to\infty}\frac{c(\gamma^{n},\gamma^{+})}{n}=\lim_{n\to\infty}\frac{d(o,\gamma^{n}o)}{n}=\ell_{d}(\gamma).$
It follows from equation (2.8), $|c(\gamma,\gamma^{+})-\ell_{d}(\gamma)|\leq 8\delta$.
For $c(\gamma,\gamma^{-})$, notice that by Proposition 2.6,
$|c(\gamma,\gamma^{+})+c(\gamma,\gamma^{-})|\leq 4\delta.$
We have $c(\gamma,\gamma^{-})+\ell_{d}(\gamma)\leq 12\delta$.
Applying this equation to $\gamma^{n}$, we have
$|c(\gamma^{n},\gamma^{\pm})\mp\ell_{d}(\gamma^{n})|=n|c(\gamma,\gamma^{\pm})\mp\ell_{d}(\gamma)|\leq 12\delta.$
Therefore, $c(\gamma,\gamma^{\pm})=\pm\ell_{d}(\gamma)$. This completes the proof.
2.D. Barycenter method and relatively Busemann function
Following Bourdon [bourdon1996birapport], Biswas [biswas2015moebius] showed that any Moebius homeomorphism $f:\partial X\to\partial Y$ between boundaries of proper, geodesically complete CAT(-1) spaces $X$, $Y$ extends to a $(1,\log 2)$-quasi-isometry $F:X\to Y$. Biswas  further generalized this result [biswas2017hyperbolic, biswas2021quasi] in a more comprehensive framework. We outline some key ideas here; for a detailed exposition, refer to [biswas2021quasi].
Consider a compact metrizable space $Z$, where an antipodal function is defined as a continuous function $\phi:Z\times Z\to[0,1]$ which satisfies symmetry, positivity, and the antipodal property where every $\xi\in Z$ has an antipode $\eta$ such that $\phi(\xi,\eta)=1$. An antipodal space is a compact Hausdorff space $Z$ with an antipodal function $\phi$. Notably, the boundary of a proper, geodesically complete, boundary-continuous Gromov hyperbolic space equipped with a visual quasi-metric serves as an example of an antipodal space.
Given an antipodal space $(Z,\phi)$, define the corresponding Moebius space $M(Z)$ to be the set of Moebius antipodal functions on $Z$, i.e., antipodal functions share the same cross-ratio as the that of the given antipodal function on $Z$. A metric is defined on $M(Z)$ by
$d_{M}(\phi_{1},\phi_{2})=\left\lVert\log\frac{d\phi_{1}}{d\phi_{2}}\right\rVert_{\infty},$
where $\frac{d\phi_{1}}{d\phi_{2}}$ is the derivative of $\phi_{1}$ with respect to $\phi_{2}$, see [biswas2021quasi]*Lemma 2.2.
One notable result from [biswas2021quasi] asserts that for a proper, geodesically complete, boundary-continuous Gromov hyperbolic space $X$, $X$ an be isometrically embedded into the space $(M(\partial X),d_{M})$ as a cobounded subset.
We call $\log\frac{d\phi_{1}}{d\phi_{2}}$ the relatively Busemann function as it represents the difference in the Busemann function when both metrics are negatively curved Riemannian. This concept is then generalized to a rough geodesic setting, and it agrees up to bounded error to the definition before.
We will have a similar construction in Section 5 based on the relatively Busemann function.
3. Main idea: length relation
In this section, we present the central method employed in this paper, focusing on the relation of stable length and the Busemann function. Before delving into the details, we find it beneficial to recall a pertinent lemma.
Lemma 3.1.
[hao2022marked]*Lemma 3.6 Let $\Gamma$ be a group. Consider an isometric action of $\Gamma$ on a non-trivial $\delta$-Gromov hyperbolic space $(X,d)$. Let $\gamma$, $\eta\in\Gamma$ with $\gamma$ acts hyperbolic on $X$. Then, for all but possibly one $n$, $\eta\gamma^{n}$ act hyperbolicly on $X$. Furthermore, if $\eta\gamma^{+}\neq\gamma^{-}$, then we have:
$\lim_{n\to\infty}(\eta\gamma^{n})^{+}=\eta\gamma^{+},$
$\lim_{n\to-\infty}(\eta\gamma^{n})^{-}=\gamma^{-}.$
With this lemma in mind, we proceed to outline the primary relation.
Proposition 3.2.
Let $\Gamma$ be a group. Given an isometric action of $\Gamma$ on a non-trivial $\delta$-Gromov hyperbolic space $(X,d)$ which is dense at infinity. Let $\gamma\in\Gamma$ act hyperbolicly on $X$, and $c$ be the rough Busemann cocycle for the action. For any $\eta\in\Gamma$ such that $\eta\gamma^{+}\neq\gamma^{-}$, we have,
$\limsup_{n\to\infty}|c(\eta,\gamma^{+})+n\ell_{d}(\gamma)-\ell_{d}(\eta\gamma^{n})|\leq|<\gamma^{+},\gamma^{-}>_{d}-<\eta\gamma^{+},\gamma^{-}>_{d}|+12\delta$
Proof.
From section 2.C, we have the following properties:
  1. (1)
    $c(\gamma,\gamma^{\pm})=\pm\ell_{d}(\gamma).$ (Proposition 2.7)
  2. (2)
    $|<x,y>_{d}-<\zeta x,\zeta y>_{d}-c(\zeta,x)-c(\zeta,y)|\leq 4\delta.$ (Proposition 2.6)
for all $x,y\in\partial X$, $\zeta\in\Gamma$ and $\gamma\in\Gamma$ that acts hyperbolicly on $X$.
Choose any $\gamma\in\Gamma$ acts hyperbolicly on $X$ with fixed points $\gamma^{+}$ and $\gamma^{-}$. Let $\eta\in\Gamma$ be any element such that $\eta\gamma^{+}\neq\gamma^{-}$. By Proposition 2.5, we have:
(3.1) $|c(\eta\gamma^{n},\gamma^{+})-c(\eta,\gamma^{n}\gamma^{+})-c(\gamma^{n},\gamma^{+})|\leq 4\delta$
for all $n$. Hence, in view of Proposition 2.7,
(3.2) $|c(\eta\gamma^{n},\gamma^{+})-c(\eta,\gamma^{+})-n\ell_{d}(\gamma)|\leq 4\delta.$
Now, for sufficiently large $n$, $\eta\gamma^{n}$ acts hyperbolicly according to Lemma 3.1. Hence, we have:
(3.3) $c(\eta\gamma^{n},\gamma^{+})-\ell_{d}(\eta\gamma^{n})=c(\eta\gamma^{n},\gamma^{+})+c(\eta\gamma^{n},(\eta\gamma^{n})^{-}.$
By Proposition 2.6, we have:
(3.4) $|<\gamma^{+},(\eta\gamma^{n})^{-}>_{d}-<\eta\gamma^{n}\gamma^{+},\eta\gamma^{n}(\eta\gamma^{n})^{-}>_{d}-c(\eta\gamma^{n},\gamma^{+})-c(\eta\gamma^{n},(\eta\gamma^{n})^{-})|\leq 4\delta.$
Note that $<\eta\gamma^{n}\gamma^{+},\eta\gamma^{n}(\eta\gamma^{n})^{-}>_{d}=<\eta\gamma^{+},(\eta\gamma^{n})^{-}>_{d}$. Taking the limit as $n\rightarrow+\infty$, using Lemma 3.1, equations (3.3) and (3.4) and (2.3), we obtain:
(3.5) $\limsup_{n\to\infty}|c(\eta,\gamma^{+})+n\ell_{d}(\gamma)-\ell_{d}(\eta\gamma^{n})-<\gamma^{+},\gamma^{-}>_{d}+<\eta\gamma^{+},\gamma^{-}>_{d}|\leq 12\delta.$
The results follow from the triangle inequality. ∎
Now, we present the first application of the established relation.
Let $\Gamma$ be a group, and $(X,d)$ a nontrivial $\delta$-Gromov hyperbolic space with a $\Gamma$-action by isometries, which is dense at infinity. According to [gromov1987hyperbolic]*8.2.D and 8.2.E, $\Gamma$ is either quasi-parabolic or non-elementary. In either case, $\Gamma$ contains a hyperbolic element $\gamma$. Without loss of generality, we may assume that the attracted fixed point $\gamma^{+}$ is a global fixed point. Let $P(N)$ be the set of all increasing sequences of positive integers, and $\phi:\Gamma\to P(N)$ be any map. Denote the image of $\xi$ by $\phi_{\xi}$ for all $\xi\in\Gamma$.
Case 1: If $\Gamma$ is quasi-parabolic, then $\gamma^{+}$ is a $\Gamma$-fixed point. Let
$E_{\phi}=\{\gamma^{-\phi(\eta)(n)}\eta^{-1}|\eta\in\Gamma,n\in\mathbb{N}^{+}\}.$
Case 2: If $\Gamma$ is non elementary, choose $\theta\in\Gamma$ such that $\theta\gamma^{-}\neq\gamma^{-}$. Let
$E^{1}_{\phi}=\{\gamma^{-\phi(\eta)(n)}\eta^{-1}|\eta\in\Gamma,n\in\mathbb{N}^{+},<\eta\gamma^{+},\gamma^{-}>_{d}\leq<\theta\gamma^{-},\gamma^{-}>_{d}+2\delta\},$
$E_{\phi}^{2}=\{\theta\gamma^{-\phi(\eta)(n)}\eta^{-1}|\eta\in\Gamma,n\in\mathbb{N}^{+},<\eta\gamma^{+},\gamma^{-}>_{d}\geq<\theta\gamma^{-},\gamma^{-}>_{d}+2\delta\}.$
And denote $E_{\phi}=E_{\phi}^{1}\cup E_{\phi}^{2}$.
We have the following:
Lemma 3.3.
The set $E_{\phi}$ is spectrally rigid.
Proof.
Let $(Y,d^{\prime})$ be any nontrivial $\delta$-Gromov hyperbolic space with a comparable $\Gamma$ action. Assume the marked length spectrum agrees on $E_{\phi}$ for these two actions. We identify $\partial X$ and $\partial Y$ by the homeomorphism $f$. Denote the Busemann cocycle for $X$ and $Y$ by $c_{1}$ and $c_{2}$.
Case 1: $\Gamma$ is quasi-parabolic. For any $\eta\in\Gamma$, applying Proposition 3.2 for $\eta$ and $\gamma$, we have:
$\limsup_{n\to\infty}|c_{1}(\eta,\gamma^{+})+\phi_{\eta}(n)\ell_{d}(\gamma)-\ell_{d}(\gamma^{-\phi_{\eta}(n)}\eta^{-1})|\\ \leq 14\delta,$
and
$\limsup_{n\to\infty}|c_{2}(\eta,\gamma^{+})+\phi_{\eta}(n)\ell_{d^{\prime}}(\gamma)-\ell_{d^{\prime}}(\gamma^{-\phi_{\eta}(n)}\eta^{-1})|\leq 14\delta.$
Here we use the fact $\ell(\zeta)=\ell(\zeta^{-1})$ for all metrics. Taking the difference, and note that $\ell_{d}(\gamma^{-\phi_{\eta}(n)}\eta^{-1}))=\ell_{d^{\prime}}(\gamma^{-\phi_{\eta}(n)}\eta^{-1})$ by assumption, we have:
lim $\limsup|c_{1}(\eta,\gamma^{+})-c_{2}(\eta,\gamma^{+})+\phi_{\eta}(n)(\ell_{d}(\gamma)-\ell_{d^{\prime}}(\gamma))|\leq 28\delta.$
Hence $\ell_{d}(\gamma)=\ell_{d^{\prime}}(\gamma)$ and $c_{1}(\eta,\gamma^{+})-c_{2}(\eta,\gamma^{+})$ are bounded. Since $\gamma^{+}$ is a fixed point of $\eta$, we have for any $n\in\mathbb{Z}$, $c_{1}(\eta^{n},\gamma^{+})-c_{2}(\eta^{n},\gamma^{+})=n(\ell_{d}(\eta)-\ell_{d^{\prime}}(\eta))$ is bounded. Therefore, $\ell_{d}(\eta)=\ell_{d^{\prime}}(\eta).$
Case 2: $\Gamma$ is non-elementary. By the definition of topology on $\partial X$ in term of Gromov product, if $<\eta\gamma^{+},\gamma^{-}>_{d}\leq<\theta\gamma^{-},\gamma^{-}>_{d}+2\delta$, then the follows is true: $<\eta\gamma^{+},\gamma^{-}>_{d^{\prime}}\leq<\theta\gamma^{-},\gamma^{-}>_{d^{\prime}}+M_{1}$ for some $M_{1}>0$.
First, let us consider $\eta\in\Gamma$ with $<\eta\gamma^{+},\gamma^{-}>_{d}\leq<\theta\gamma^{-},\gamma^{-}>_{d}+2\delta$. Applying Proposition 3.2 for $\eta$ and $\gamma$, we have
$\limsup_{n\to\infty}|c_{1}(\eta,\gamma^{+})+\phi_{\eta}(n)\ell_{d}(\gamma)-\ell_{d}(\gamma^{-\phi_{\eta}(n)}\eta^{-1})|\\ \leq<\gamma^{+},\gamma^{-}>_{d}+<\eta\gamma^{+},\gamma^{-}>_{d}+14\delta.$
and similarly
$\limsup_{n\to\infty}|c_{2}(\eta,\gamma^{+})+\phi_{\eta}(n)\ell_{d^{\prime}}(\gamma)-\ell_{d^{\prime}}(\gamma^{-\phi_{\eta}(n)}\eta^{-1})|\leq A+12\delta+M_{1},$
where $A=<\gamma^{+},\gamma^{-}>_{d^{\prime}}+<\theta\gamma^{+},\gamma^{-}>_{d^{\prime}}$. By assumption,
$\ell_{d}(\gamma^{-\phi_{\eta}(n)}\eta^{-1})=\ell_{d^{\prime}}(\gamma^{-\phi_{\eta}(n)}\eta^{-1}).$
Hence we have there exist a constant $M_{\gamma,\theta,X,Y}$,
lim $\limsup|c_{1}(\eta,\gamma^{+})-c_{2}(\eta,\gamma^{+})+\phi_{\eta}(n)(\ell_{d}(\gamma)-\ell_{d^{\prime}}(\gamma))|\leq M_{\gamma,\theta,X,Y}.$
It follows that $\ell_{d}(\gamma)-\ell_{d^{\prime}}(\gamma)=0$, and $|c_{1}(\eta,\gamma^{+})-c_{2}(\eta,\gamma^{+})|$ is uniformly bounded for $\eta$.
Second, let us consider $\eta\in\Gamma$ with $<\eta\gamma^{+},\gamma^{-}>_{d}\geq<\theta\gamma^{-},\gamma^{-}>_{d}+2\delta$. By hyperbolic inequality (2.2),
$<\theta\gamma^{-},\gamma^{-}>_{d}\geq\min\{<\eta\gamma^{+},\theta\gamma^{-}>_{d},<\eta\gamma^{+},\gamma^{-}>_{d}\}-2\delta.$
Hence, $<\eta\gamma^{+},\theta\gamma^{-}>_{d}\leq<\theta\gamma^{-},\gamma^{-}>_{d}+2\delta.$ And in this case, there exist $M_{2}>0$ so that $<\eta\gamma^{+},\theta\gamma^{-}>_{d^{\prime}}\leq<\theta\gamma^{-},\gamma^{-}>_{d^{\prime}}+M_{2}.$ By applying the first case to the pair $\eta\theta^{-1}$ and $\theta\gamma\theta^{-1}$, we know that $|c_{1}(\eta\theta^{-1},\theta\gamma^{+})-c_{2}(\eta\theta^{-1},\theta\gamma^{+})|$ is uniformly bounded.
By the cocycle identity (2.5),
$|c_{i}(\eta,\gamma^{+})|\leq|c_{i}(\eta\theta^{-1},\theta\gamma^{+})-c_{i}(\theta^{-1},\theta\gamma^{+})|+8\delta.$
Therefore, there exists $K>0$ such that $|c_{1}(\eta,\gamma^{+})-c_{2}(\eta,\gamma^{+})|\leq K$ for all $\eta\in\Gamma$.
Now, for any $\zeta$ and $\xi\in\Gamma$, the cocycle identity implies that
$|c_{i}(\zeta,\xi\gamma^{+})|\leq|c_{i}(\zeta\xi,\gamma)-c_{i}(\xi,\gamma^{+})|+8\delta.$
Hence $|c_{1}(\zeta,\xi\gamma^{+})-c_{2}(\zeta,\xi\gamma^{+})|\leq 2K+16\delta.$ By Proposition 2.4 and the fact that $\Gamma\gamma^{+}$ is dense, we have, for all hyperbolic element $\eta\in\Gamma$, we have $|c_{1}(\eta,\eta^{+})-c_{2}(\eta,\eta^{+})|\leq 2K+24\delta.$ Applying this equation to $\eta^{n}$, by Proposition 2.7, we have $\ell_{d}(\eta)=\ell_{d^{\prime}}(\eta)$ for all hyperbolic element $\eta.$
This completes the proof.
4. Proof of Theorem D and E
4.A. Proof of Theorem D
The proof of Theorem D is an application of Lemma 3.3. First, let us restate Theorem D.
Theorem 4.1.
Let $f:\mathbb{R}\to\mathbb{R}^{+}$ such that $\lim_{T\to\infty}f(T)=\infty$ and $\Gamma$ be a group. There exists a subset $E\subset\Gamma$ so that: given two comparable actions of $\Gamma$ on two non-trivial Gromov hyperbolic spaces $(X,d)$ and $(Y,d^{\prime})$ which are dense at infinity, if $\ell_{d}(\gamma)=\ell_{d^{\prime}}(\gamma)$ for all $\gamma\in E$, then $\ell_{d}=\ell_{d^{\prime}}$ on $\Gamma$. Furthermore, we can choose $E$ so that:
  1. (1)
    the growth of $E$ with respect to $d$ (or/and $d^{\prime}$) satisfying:
    #$\#\{\gamma\in E|\ell_{d}(\gamma)\leq T\}\leq f(T).$
  2. (2)
    If the action is quasi-parabolic, for ant $\gamma\in\Gamma$ acts on $X$ hyperbolically, the limit set of $E$ contains only the $\gamma$-fixed point which is not $\Gamma$-fixed.
  3. (3)
    If the action is non-elementary, for any two given $\gamma$ and $\theta\in\Gamma$ with $\gamma$ acts on $X$ hyperbolically and $\theta\gamma^{-}\neq\gamma^{-}$, the limit set of $E$ contains at most two points: $\gamma^{-}$ and $\theta\gamma^{-}$.
Proof.
Fix a base point $o\in X$.
Case 1: $\Gamma$ is quasi-parabolic Let $\Gamma$ be a fixed hyperbolic element such that $\gamma^{+}$ is a $\Gamma$-fixed point. Note that by Lemma 3.1, for any $\eta\in\Gamma$, for $n$ big enough, $\gamma^{-n}\eta^{-1}$ is hyperbolic, and $|\ell_{d}(\gamma^{-n}\eta^{-1})-n\ell_{\gamma}|\leq K_{\eta,\gamma}$ by Proposition 3.2. Enumrate the element of $\Gamma$ by $\eta_{i}$, $i\in\mathbb{N}$. For each $i\in\mathbb{N}$, choose a sequence $n^{i}$ so that
  1. (1)
    #$\#\{m|\ell_{d}(\gamma^{-n^{i}(m)}\eta^{-1})\leq T\}\leq\frac{f(T)}{2^{i}}$.
  2. (2)
    $<\gamma^{-n^{i}(m)}\eta^{-1}o,\gamma^{-}>_{d}\geq 1000i^{2}(\delta+1)$ for all $m>1$.
Let $\phi$ be the map $\phi(\eta_{i})=n^{i}$, and set $E=E_{\phi}.$ Then the only limit point of $E$ is $\gamma^{-}$, and $E$ is spectrally rigid by Lemma 3.3. By the construction, the growth rate of $E$ is less that $f$.
Case 2: $\Gamma$ is non-elementary. Let $\gamma$ be a hyperbolic element, and $\theta\in\Gamma$ with $\theta\gamma^{-}\neq\gamma^{-}$. Note that by Lemma 3.1, for any $\eta\in\Gamma$ with $\eta\gamma^{+}\neq\gamma^{-}$ , for $n$ big enough, $\gamma^{-n}\eta^{-1}$ is hyperbolic, and $|\ell_{d}(\gamma^{-n}\eta^{-1})-n\ell_{\gamma}|\leq K_{\eta,\gamma}$ by Proposition 3.2. Enumerate the element such that $<\eta\gamma^{+},\gamma^{-}>_{d}\leq<\theta\gamma^{-},\gamma^{-}>_{d}+2\delta$ by $\eta_{2n-1}$ for $n\geq 1$ and other elements by $\eta_{2n}$ for $n\geq 1$.
For each $i=2k+1$, choose a sequence $n^{i}$ so that
  1. (1)
    #$\#\{m|\ell_{d}(\gamma^{-n^{i}(m)}\eta^{-1})\leq T\}\leq\frac{f(T)}{2^{i}}$.
  2. (2)
    $<\gamma^{-n^{i}(m)}\eta^{-1}o,\gamma^{-}>_{d}\geq 1000i^{2}(\delta+1)$ for all $m>1$.
For each $i=2k$, choose a sequence $n^{i}$ so that
  1. (1)
    #$\#\{m|\ell_{d}(\theta\gamma^{-n^{i}(m)}\eta^{-1})\leq T\}\leq\frac{f(T)}{2^{i}}$.
  2. (2)
    $<\theta\gamma^{-n^{i}(m)}\eta^{-1}o,\theta\gamma^{-}>_{d}\geq 1000i^{2}(\delta+1)$ for all $m>1$.
Define a map $\phi:\Gamma\to P(n)$ by $\phi(\eta_{i})=n^{i}$. The $E_{\phi}$ is spectrally rigid by Lemma 3.3, with a growth rate less than $f$, and has two limit points $\gamma^{-}$ and $\theta\gamma^{-}$. ∎
4.B. Proof of Theorem E
The proof of Theorem E is similar to the proof of Theorem D.
Let $\Gamma$ be a trosion-free word hyperbolic group, and consider a faithful action of $\Gamma$ on a nontrivial Gromov hyperbolic space that is dense at infinity. First, we construct the set $E^{\prime}_{\phi}$ for a map $\phi:\Gamma\to P(N)$. Let $\gamma\in\Gamma$ act as a hyperbolic isometry on $X$. Without loss of generality, assume $\gamma^{-}$ is not a $\Gamma$-fixed point, and has a dense orbit. Let $E(\gamma)$ be the maximal elementary subgroup of $\Gamma$ contains $\gamma$, (see [minasyan2010normal]*Lemma 2.4). there is an element $\eta_{\gamma}\in\Gamma\setminus E(\gamma)$ so that $\eta_{\gamma}\gamma^{-}\neq\gamma^{+}$. This is true because, by (2) of Lemma 2.4 in [minasyan2010normal], for all $\xi\in E(\gamma)$, $\xi(\gamma^{-})$ is one of the two points $\gamma^{\pm}$ and the orbit of $\gamma^{-}$ is dense. Define the set $E^{\prime}_{\phi}$ as follows:
$E^{\prime}_{\phi}:=\{\gamma^{-\phi_{\gamma}(n)}\eta_{\gamma}^{-1}|\gamma\in\Gamma\text{ hyperbolic, }n\in\mathbb{N}^{+}\}.$
Lemma 4.2.
The set $E^{\prime}_{\phi}$ detects the ratio of the marked length spectrum.
Proof.
Let $(Y,d^{\prime})$ be any nontrivial $\delta$-Gromov hyperbolic space with a comparable $\Gamma$ action. Assume there exists $C>1$ such that the marked length spectrum satisfies $\frac{1}{C}\leq\frac{\ell_{d}(\zeta)}{\ell_{d^{\prime}}(\zeta)}\leq C$ for all $\zeta\in E^{\prime}_{\phi}$ We identify $\partial X$ and $\partial Y$ by the homeomorphism $f$. Denote the Busemann cocycle for $X$ and $Y$ by $c_{1}$ and $c_{2}$.
Let $\gamma$ be a hyperbolic element. Applying Proposition 3.2 for $\eta_{\gamma}$ and $\gamma$, we have
$\limsup_{n\to\infty}|c_{1}(\eta_{\gamma},\gamma^{+})+\phi_{\gamma}(n)\ell_{d}(\gamma)-\ell_{d}(\eta_{\gamma}\gamma^{\phi_{\gamma}(n)})|\leq<\gamma^{+},\gamma^{-}>_{d}+<\eta_{\gamma}\gamma^{-},\gamma^{-}>_{d}+12\delta,$
$\limsup_{n\to\infty}|c_{2}(\eta_{\gamma},\gamma^{+})+\phi_{\gamma}(n)\ell_{d^{\prime}}(\gamma)-\ell_{d^{\prime}}(\eta_{\gamma}\gamma^{\phi_{\gamma}(n)})|\leq<\gamma^{+},\gamma^{-}>_{d^{\prime}}+<\eta_{\gamma}\gamma^{-},\gamma^{-}>_{d^{\prime}}+12\delta,$
Therefore, $\phi_{\gamma}(n)\ell_{d}=\ell_{d}(\eta_{\gamma}\gamma^{\phi_{\gamma}(n)})+M_{n,\gamma}$ and $\phi_{\gamma}(n)\ell_{d^{\prime}}=\ell_{d^{\prime}}(\eta_{\gamma}\gamma^{\phi_{\gamma}(n)})+M^{\prime}_{n,\gamma}$, where both $M_{n,\gamma}$ and $M^{\prime}_{n,\gamma}$ is bounded. Hence
(4.1) $\frac{\ell_{d}(\gamma)}{\ell_{d^{\prime}}(\gamma)}=\frac{\ell_{d}(\eta_{\gamma}\gamma^{\phi_{\gamma}(n)})+M_{n,\gamma}}{\ell_{d^{\prime}}(\eta_{\gamma}\gamma^{\phi_{\gamma}(n)})+M^{\prime}_{n,\gamma}}$
Since $\gamma$ is hyperbolic, $\phi_{\gamma}(n)\ell_{d}=\ell_{d}(\eta_{\gamma}\gamma^{\phi_{\gamma}(n)})+M_{n,\gamma}$ implies
$\lim_{n\to\infty}\ell_{d}(\eta_{\gamma}\gamma^{\phi_{\gamma}(n)})=\infty.$
The same is true for the $d^{\prime}$ metric on $Y$. Taking limit of equation (4.1), we have
$\frac{1}{C}\leq\frac{\ell_{d}(\gamma)}{\ell_{d^{\prime}}(\gamma)}\leq C.$
This completes the proof. ∎
Proof of Theorem E.
Same as in the proof of Theorem D, by taking $\phi_{\gamma}$ big and sparse enough, we have a sparse subset. The only thing left is to show that $E^{\prime}_{\phi}$ contains only primary elements. Indeed, without loss of generality, we assume $\gamma$ is a primary element in $\Gamma$. This follows from Lemma 3.6 from [minasyan2010normal]. ∎
Remark 4.3.
Note that when $\Gamma$ is quasi-parabolic, and by choosing $\phi_{\gamma}$ big enough, the set $E^{\prime}_{\phi}$ has only one limit point, which is the $\Gamma$ fixed point.
5. Rough cross-ratio of roughly geodesic hyperbolic space and Busemann functions
In this section, we prove Theorem 5.1, which is essentially the first part of the proof of Theorem B.
Theorem 5.1.
Let $(X,d)$, $(Y,d^{\prime})$ be two comparable nontrivial $\Gamma$-Gromov hyperbolic spaces such that $\Gamma$ is dense at infinity for $X$ and $Y$. Let $<,>_{d}$ and $<,>_{d^{\prime}}$ be the Gromov products on $\partial X$ and $\partial Y$ for some base points, respectively. If $\ell_{d}=\ell_{d^{\prime}}$, then there exists $L>0$ such that
$|<x,y>_{d}-<f(x),f(y)>_{d^{\prime}}|\leq L$
for all $x\neq y\in\partial X$.
The theorem is known before for cocompact actions. See Furman [F].
Proof of theorem 5.1.
Since the two actions are comparable, there is a continuous homeomorphism $f:\partial X\to\partial Y$. We identify $\partial X$ and $\partial Y$ via $f$. Let $c_{1}$ and $c_{2}$ denote the Busemann cocycle of $X$ and $Y$, respectively.
Let $\bar{c}=c_{1}-c_{2}$ and $g=<x,y>_{d}-<x,y>_{d^{\prime}}$. The first step of the proof of Lemma 3.3 shows that $\bar{c}$ is bounded.
By Proposition 2.6, $dg$ is a bounded cocycle. Here, $dg(\gamma,(x,y))=g(x,y)-g(\gamma x,\gamma y)$. Hence, there exists a bounded function $h$ such that $dg=dh$. Indeed, we may take $h(x)=\sup_{\gamma\in\Gamma}dg(\gamma,x)$. Since $g-h$ is $\gamma$-invariant, $g-h$ is constant on any $\Gamma$-orbit. There exists a dense orbit. In view of equation (2.3), $g-h$ is bounded. Hence, $g$ is bounded.
This completes the proof. ∎
6. Infinite-barycenter map
We conclude the proof of Theorem B by presenting the hyperbolic filling of a boundary.
Consider a roughly hyperbolic manifold $(X,d)$ with a designated base point $o$. For any real numbers $K>100$ and $\delta>0$, let $\mathcal{D}^{R}_{\delta,K}$ be the collection of pointed, $(K,\delta)$-roughly hyperbolic manifold metrics $D$ on $X$, such that for all $(x,y,z,w)\in\partial^{(4)}(X)$:
  1. (1)
    $|<x,y>_{d}-<x,y>_{D}|<\infty.$
  2. (2)
    $\sup_{y\in\partial X}\inf_{x\in\partial X}\{<x,y>_{D}\}\leq K-1.$
  3. (3)
    $|(x,y;z,w)_{d}-(x,y;z,w)_{D}|\leq K,$
where $(x,y;z,w)$ denotes the rough cross-ratio (2.4). The base point of $D$ is denoted by $o_{D}$, and the Gromov product for $D$ is understood with respect to the base point $o_{D}$.
Remark 6.1.
The space $\mathcal{D}^{R}_{\delta,K}$ is similar to [oregon2023space]*Defnition 1.1. $\mathcal{D}^{R}_{\delta,K}$ is a point in $\mathcal{D}(\Gamma)$ with further restrictions on the geometries since all the spaces are roughly isometric. And the metric we defined here is essentially to find the best $A$ in [oregon2023space]*Defnition 1.2.
We introduce a metric $\rho$ on $\mathcal{D}^{R}_{\delta,K}$ defined by
$\rho(D_{1},D_{2})=\sup_{x\neq y\in\partial X}\{|<x,y>_{D_{1}}-<x,y>_{D_{2}}|\}.$
For every point in $p\in X$, there exists an element $D_{p}\in\mathcal{D}^{R}_{\delta,K}$ for sufficiently large $K$. This element is the same metric space $(X,d)$ with base point $p$. It is evident from the definition and the triangle inequality that $\rho(D_{p},D_{q})\leq d(p,q)$. We denote this embedding map by $i_{X}$.
Firstly, we establish that $i_{X}$ is a roughly geodesic embedding. Secondly, we show that the image of $i_{X}$ is cobounded.
6.A. The embedding is roughly isometric embedding
Lemma 6.2.
The embedding $i_{X}$ is a roughly geodesic embedding.
Proof.
Consider any $p$, $q\in X$. Let $r$ be a $K$-roughly geodesic with $r(0)=p$, $r(t)=q$, $r(\infty)=x\in\partial X$ such that $|t-d(p,q)|\leq K$. Choose $y\in\partial X$ such that $<x,y>_{p}\geq d(p,q)$. Direct computation and equation (2.1) yeild:
$|(<x,y>_{p}-<x,y>_{q})-(<y,q>_{p}-<x,p>_{q})|\leq 8\delta.$
As $r$ is roughly geodesic, applying equation (2.5) and (2.1) gives:
$<x,p>_{q}\leq\frac{3}{2}K+2\delta.$
Taking a limit of the triangle inequality, we have $<x,q>_{p}\leq d(p,q)$. By the hyperbolic inequality (2.2), equation (2.6) and (2.3),
$<y,q>_{p}\geq\min\{<y,x>_{p},<x,q>_{p}\}-2\delta=<x,q>_{p}-\delta\geq d(p,q)-\frac{5}{2}K-3\delta.$
Thus:
$\rho(D_{p},D_{q})\geq|<x,y>_{p}-<x,y>_{q}|\geq d(p.q)-4K-13\delta.$
On the other hand, for any $x\neq y\in\partial X$, consider sequences $(r_{i})$ and $(s_{j})$ such that:
$\lim_{i,j}<r_{i},s_{j}>=<x,y>_{p}$
with $r_{i}\to x$, $s_{j}\to y$, taking the limit using this sequence, and equation (2.3) along with the triangle inequality, we have
$|<x,y>_{p}-<x,y>_{q}|\leq d(p,q)+4\delta.$
Hence
$\rho(D_{p},D_{q})\leq d(p,q)+4\delta.$
This completes the proof. ∎
6.B. Coboundedness
First, we introduce the rough version of the relatively Busemann function. Given $D_{1}$, $D_{2}\in\mathcal{D}^{R}_{\delta,K}$, let
$f_{D_{1},D_{2}}(x)=\frac{1}{2}\liminf_{y\to x}<x,y>_{D_{1}}-<x,y>_{D_{2}}.$
Then we have $<x,y>_{d_{1}}-<x,y>_{d_{2}}-f_{D_{1},D_{2}}(x)-f_{D_{1},D_{2}}(y)$ is the same as
$\frac{1}{2}\liminf_{z\to x,w\to y}[(x,y;z,w)_{D_{1}}-(x,y;z,w)_{D_{2}}+T(z,w)]$
where $T(z,w)=<x,y>_{D_{1}}-<z,w>_{D_{1}}-<x,y>_{D_{2}}+<z,w>_{D_{2}}$. By equation (2.3), lim $\liminf|T(z,w)|\leq 8\delta$. Therefore
(6.1) $|<x,y>_{D_{1}}-<x,y>_{D_{2}}-f_{D_{1},D_{2}}(x)-f_{D_{1},D_{2}}(y)|\leq\frac{1}{2}K+4\delta.$
We call $f_{D_{1},D_{2}}$ the relatively Busemann function of $D_{1}$ and $D_{2}$. Here are a few properties of the relatively Busemann function:
Lemma 6.3.
The supremum and infimum of the Busemann function are roughly opposite.
Proof.
Let $x\in X$ be such that $f_{D_{1},D_{2}}(x)\geq\sup f_{D_{1},D_{2}}-\delta$. Then, by definition, there exist $x\neq y\in\partial X$, such that $<x,y>_{D_{1}}\leq K$, in view of Equation (6.1),
$-4\delta-\frac{1}{2}K+f_{D_{1},D_{2}}(x)+f_{D_{1},D_{2}}(y)\leq<x,y>_{D_{1}}-<x,y>_{D_{2}}\leq<x,y>_{D_{1}}\leq K.$
Hence $\inf(f_{D_{1},D_{2}})\leq-\sup f_{D_{1},D_{2}}+5\delta+\frac{3}{2}K$. A similar consideration by letting $x\neq y$ such that $f_{D_{1},D_{2}}(y)\leq\inf f_{D_{1},D_{2}}+\delta$, and $<x,y>_{D_{2}}\leq K$ leads to
$4\delta+\frac{1}{2}K+f_{D_{1},D_{2}}(x)+f_{D_{1},D_{2}}(y)\geq<x,y>_{D_{1}}-<x,y>_{D_{2}}\geq-<x,y>_{D_{2}}\geq-K.$
Hence $\inf f_{D_{1},D_{2}}\geq-\sup f_{D_{1},D_{2}}-\frac{3}{2}K-5\delta$. Therefore,
$||\inf f_{D_{1},D_{2}}+\sup f_{D_{1},D_{2}}||\leq 5\delta+\frac{3}{2}K.$
Lemma 6.4.
$\rho(D_{1},D_{2})$ is roughly equal to $2\sup f_{D_{1},D_{2}}$.
Proof.
By definition of $\rho$ and equation (6.1), it is clear that
$\rho(D_{1},D_{2})\leq 2\max\{|\inf{|f_{D_{1},D_{2}}|}|,|\sup{f_{D_{1}.D_{2}}}|\}+\frac{1}{2}K+4\delta.$
Hence, by Lemma 6.3,
(6.2) $\rho(D_{1},D_{2})\leq 2|\sup{f_{D_{1}.D_{2}}}|+\frac{7}{2}K+14\delta.$
On the other hand, consider a sequence $w_{n}\in\partial X$ converge to $x$, by (6.1),
$\begin{array}[]{cl}|f_{D_{1},D_{2}}(x)-f_{D_{1},D_{2}}(w_{n})|\leq&\frac{1}{2}K+4\delta+|<x,y>_{D_{1}}-<w_{n},y>_{D_{1}}|\\ &+|<x,y>_{D_{2}}-<w_{n},y>_{D_{2}}|.\end{array}$
For $n$ big enough, according to equation (2.3), we have
$|f_{D_{1},D_{2}}(x)-f_{D_{1},D_{2}}(w_{n})|\leq\frac{1}{2}K+8\delta.$
Hence $\rho(D_{1},D_{2})\geq|<x,w_{n}>_{D_{1}}-<x,w_{n}>_{D_{2}}|\geq 2f_{D_{1},D_{2}}(x)-K-12\delta$ for all $x$. In particular,
(6.3) $\rho(D_{1},D_{2})\geq 2|\sup f_{D_{1}.D_{2}}|-K-12\delta.$
Now we are ready to show that the image is cobounded.
Lemma 6.5.
Let $(X,d)$ be a roughly $(\delta^{\prime},K^{\prime})$-hyperbolic manifold. Then, for $\delta$, $K$ sufficiently large, $i_{X}:X\to\mathcal{D}^{R}(\delta,K)$ is a roughly isometry. Indeed, for any point $D\in\mathcal{D}^{R}_{\delta,K}$, there is a point $p\in X$ so that $\rho(D,D_{p})\leq 1000K+1000\delta$.
Proof.
For all different points $p$, $q\in X$ and $(x,y,z,w)\in\partial^{(4)}X$. By the triangle inequality and equation (2.1), taking 4 sequence converging to $x,y,z,w$ respectively, we have
  1. (1)
    $|<x,y>_{p}-<x,y>_{q}|\leq d(p,q)+4\delta.$
  2. (2)
    $|(x,y;z,w)_{p}-<x,y;z,w>_{q}|\leq 16\delta.$
Since $X$ is $K^{\prime}$-visual, there is a $K^{\prime}$-geodesic connect any two point $p\in X$ and $x\in\partial X$. By the fact that $X$ is roughly complete, this $K^{\prime}$ rough geodesic extends to a bi-infinite $K^{\prime}$ rough geodesic. Denote the other end by $y$. By equation (2.5) and (2.1), we have:
$<x,y>_{p}\leq\frac{3}{2}K^{\prime}+2\delta.$
Let $\delta\geq\delta^{\prime}$ and $K\geq\frac{3}{2}K^{\prime}+16\delta+1$. We have $i_{X}:X\to\mathcal{D}^{R}(\delta,K)$ is a rough isometric embedding by Lemma 6.2.
Assume the image of $i_{X}$ is not $1000(K+\delta)$ dense, let $p\in X$ such that $1000K+1000\delta<\rho(D_{p},D)<\inf\{\rho(D,D_{q}|q\in X)\}+\delta$. Let
$A=\left\{{x\in\partial X}\ /\ {f_{D_{p},D}\geq\frac{\rho(D_{p},D)}{2}-100K-100\delta}\right\}.$
By Lemma 6.4, $A$ is not empty. Taking one point $x\in A$, and from the roughly geodesic $\phi(t)$ with respected to $(X,d)$ who connected $p$ and $x$ with $\phi(0)=p$. Let $q=\phi(50K+50\delta)$. Then, for any point $y\notin A$, by definition of rough geodesics, we have
$f_{D_{q},D}(y)\leq f_{D_{p},D}(y)+\frac{1}{2}d(p,q)+\delta\leq\frac{\rho(D_{p},D)}{2}-100K-100\delta+26K+26\delta.$
When $z\in A$, notice that $<x,z>_{D}\geq 0$. In view of equation (6.1),
$<x,z>_{D_{p}}\geq\rho(D_{p},D)-200K-200\delta-\frac{1}{2}K-4\delta.$
By the definition of the relatively Busemann function, we have:
$f_{D_{q},D}(z)\leq f_{D_{p},D}(z)+\frac{1}{2}\limsup_{w\to z}<z,w>_{q}-<z,w>_{p}.$
It is a direct computation and equation (2.1) that shows that:
$<z,w>_{q}-<z,w>_{p}\leq<w,p>_{q}-<z,q>_{p}+8\delta.$
First, since $x$ is the limit of a rough geodesic, equation (2.1) and
$K+2\delta\geq(x,p)_{q}\geq\min\{<w,p>_{q},<x,w>_{q}\}-2\delta,$
implies that $<w,p>_{q}\leq K+4\delta$. Second,
$<z,q>_{p}\geq\min\{(x,z)_{p},<x,q>_{p}\}-2\delta.$
Therefore, $<z,q>_{p}\geq d(p,q)-K-2\delta=48K+48\delta$ from equation (2.5). It follows that
$f_{D_{q},D}(z)\leq f_{D_{p},D}(z)-\frac{47}{2}K-18\delta.$
Therefore by equation (6.3),
$\sup_{z\in A}F_{D_{q},D}(z)\leq\frac{\rho(D_{p},D)}{2}+\frac{1}{2}K+6\delta-\frac{47}{2}K-18\delta.$
And recall for $y\notin A$,
$\sup_{y\notin A}F_{D_{q},D}(y)\leq\frac{\rho(D_{p},D)}{2}-74K-74\delta.$
Finally, according to equation (6.2),
$\rho(D_{q},D)\leq 2\sup F_{D_{q},D}+3.5K+14\delta\leq\rho(D_{p},D)-42.5K-10\delta.$
Contradict to the choice of $p$. Hence, the image of $X$ is $1000K+1000\delta$ dense.
This completes the proof. ∎
6.C. Proof of Theorem B
Proof.
Choose $\delta$, $K$ bigger enough so that $i_{X}$ and $i_{Y}$ are rough isometrics. Since both maps are $\Gamma$-equivalent. The result follows. ∎
7. Proof of Theorem C
Proof.
Since $\ell_{d}(h\gamma)=\ell_{d}(\gamma^{-1}h^{-1})$. Up to switching $h$, we may assume the marked length spectrum agrees on a left coset $g\Gamma^{\prime}$.
Same as in the proof of Theorem B, we identify $\partial X$ and $\partial Y$ via the comparable map $f$. Let $c_{1}$ and $c_{2}$ denote the Busemann cocycle of $X$ and $Y$, respectively. Let $c=c_{1}-c_{2}$ and $g(x,y)=<x,y>_{d}-<x,y>_{d^{\prime}}$. Then form section 2.C, we have:
(7.1) $c(h\gamma,(h\gamma)^{\pm})=0,$
(7.2) $|g(x,y)-g(\eta x,\eta y)-c(\eta,x)-c(\eta,y)|\leq 8\delta.$
for all $\gamma\in\Gamma^{\prime}$, $\eta\in\Gamma$, $x$, $y\in\partial X$.
First, we show that $\ell_{d}=\ell_{d^{\prime}}$ on $\Gamma^{\prime}$. By Proposition 2.7, it is equivalent to that for all hyperbolic element $\gamma\in\Gamma^{\prime}$,
$\lim_{n}\frac{c(\gamma^{n},\gamma^{-})}{n}=0.$
[…]
(7.3) $|g(\gamma^{-},(h\gamma^{n})^{+})-g(h\gamma^{-},(h\gamma^{n})^{+})-c(h\gamma^{n},\gamma^{-})-c(h\gamma^{n},(h\gamma^{n})^{+})|\leq 8\delta.$
Now we study each term in equation (7.3), and connect it to $c(\gamma^{n},\gamma^{-})$.
  1. (1)
    In view of Lemma 3.1 and equation (2.3), we have
    $\limsup_{n}|g(\gamma^{-},(h\gamma^{n})^{+})|\leq|g(\gamma^{-},h\gamma^{+})|+4\delta.$
  2. (2)
    For the same reason,
    $\limsup_{n}g(h\gamma^{-},(h\gamma^{n})^{+})\leq|g(h\gamma^{-},h\gamma^{+})|+4\delta.$
  3. (3)
    By roughly cocycle identity (2.5)
    $|c(h\gamma^{n},\gamma^{-})-c(h,\gamma^{-})-c(\gamma^{n},\gamma^{-})|\leq 4\delta.$
    Since by the triangle inequality, $c(h,\gamma^{-})\leq d(o,ho)+d^{\prime}(o^{\prime},ho^{\prime})$ where $o$ and $o^{\prime}$ are base point of $X$ and $Y$, respectively. Hence, $c(h\gamma^{n},\gamma^{-})$ is uniformly close to $c(\gamma^{n},\gamma^{-})$.
  4. (4)
    By equation 7.1,
    $c(h\gamma,(h\gamma)^{\pm})=0.$
Put them back in equation (7.3), we have, for $n$ big enough,
$|c(\gamma^{n},\gamma^{-})|\leq 20\delta+d(o,ho)+d^{\prime}(o^{\prime},ho^{\prime})+|g(\gamma^{-},h\gamma^{+})|+|g(h\gamma^{-},h\gamma^{+})|+1.$
Hence $\lim_{n}\frac{c(\gamma^{n},\gamma^{-})}{n}=0$.
Now, view $X$ and $Y$ as a $\Gamma^{\prime}$-space, by proof of Theorem B. Both $i_{X}:X\to\mathcal{D}_{\delta,K}$ and $i_{Y}:Y\to\mathcal{D}_{\delta,K}$ are rough isometics for some sufficient big $K$. Since both $\Gamma$ actions preserve the rough cross-ratio, $i_{X}$ and $i_{Y}$ are (roughly) $\Gamma$-equivariant. Hence, the result follows. ∎